The measure of an angle is 4y°. Find the measure of its supplement

Answer by jim_thompson5910(35256)

Let x = supplment of angle 4y

The two angles must add to 180 since they are supplements, so

x+4y = 180

x+4y-4y = 180-4y

x+0y = 180-4y

x+0 = 180-4y

x = 180-4y

x = -4y + 180

So the final answer is -4y + 180